OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (mrnood):

what do oyu know about the equation for a straight line y=mx+b?

OpenStudy (anonymous):

so i put it into y= mx+b and then plug in the numbers

OpenStudy (mrnood):

from the original data (slope and intercept - write the equation in y=mx+b form THEN Rearrang all the answers into the same form - one of them will be th e same as your equaiton

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

can you transform one of the equations into y +mx+b and tell me how to do it so i can do the other ones

OpenStudy (anonymous):

\[ y - 1 = 4(x - 3)\] distribute on the right \[y-1=4x-12\] add \(1\) so solve for \(y\) \[y=4x-11\] steps are always the same

OpenStudy (anonymous):

alright thank you:)